OpenStudy (anonymous):

How do you solve this:

OpenStudy (anonymous):

\[(1/r) + (r/(2-r))=-7/4\]

OpenStudy (anonymous):

\(\dfrac{1}{r}+\dfrac{r}{2-r}=-\dfrac{7}{4}\) It looks like you need a common denominator.

OpenStudy (anonymous):

right but it gets a little messy later and I got confused

OpenStudy (anonymous):

I multiplied the entire equation through by each denominator to eliminate the fractions.

OpenStudy (anonymous):

One at a time will help to reduce the confusion.

OpenStudy (anonymous):

Do the denominator calculus of the left equation. Will be: r(r - 2). After, do the normal account. Will be: \[\dfrac{-r^2-r+2}{(r-2)r}=-\dfrac{7}{4}\] Now continue the calculus doing the cross multiplying.